MODEL FOR LINE-MIXING EFFECTS IN THE INFRARED ν_{3} BAND OF CH_{4}. APPLICATION TO AT-MOSPHERIC SPECTRA

\underline{D. PIERONI}, NGUYEN-VAN-THAN, C. BRODBECK, A. VALENTIN, C. CLAVEAU, C. CAMY-PEYRET and J.-M. HARTMANN, Laboratoire de Physique Moléculaire et Applications, Bât 350. Université Paris-Sud, 91405 Orsay Cedex, France; T. GABARD and J.-P. CHAMPION, Laboratoire de Physique de l’Université de Bourgogne, Unité associée au CNRS, Faculté des Sciences Mirande, B. P. 400, 21011 Dijon Cedex, France.

Recent studies, using laboratory and atmospheric spectra, have demonstrated the influence of line-mixing effects on CH_{4} infrared absorption. This phenomenon can significantly affect remote sensing results for atmospheric applications and the contribution of CH_{4} to the green house effect. We propose a model of the relaxation operator in order to account for these effects in the computations of the absorption. The model was built using a semi-classical calculation of collisional energy transfer rates^{a} and few empirical parameters. Results obtained for N_{2}-broadened spectra of the ν_{3} band of CH_{4} (near 3 μm) including laboratory measurements covering different pressure and temperature ranges, and atmospheric spectra will be presented.

aT. Gabard, J. Quant. Spectrosc. Radiat. Transfer Vol. 57, no. 2, pp. 177-196, 1997.